# Controllable Reference-Based Real-World Remote Sensing Image Super-Resolution with Generative Diffusion Priors

### v1

@@ 5.6.2 Effect of Reference Similarity
5.6.3 Local-Global Reference Strength Control
Although our proposed CRefDiff demonstrates a strong ability to identify land cover changes between the LR and reference images, some failure cases involving incorrect reference usage still occur under certain challenging scenarios. To address this issue, we propose a Reference Strength Control strategy, which allows users to manually correct such cases. During inference, users can either supply a global scalar factor $s$ to adjust the overall reliance on reference information, or provide a spatial mask $m$ to modulate reference strength locally. As illustrated in Fig. 20, a representative example is shown. In the highlighted region, the HR and reference images exhibit significant structural differences. However, the model still incorporates some misleading reference cues, resulting in erroneous structures in the reconstructed output (as indicated by “$s=1$”). To alleviate this, the user can progressively reduce the global strength factor to suppress reference guidance. Alternatively, a binary mask specifying unreliable regions can be provided to locally weaken the influence of reference features in those areas.
@@ 5.6.4 Model Efficiency
To compare the runtime efficiency of different methods, we evaluate the inference time and GPU memory consumption for a 480$\times$480 image on an NVIDIA GeForce RTX 3090 GPU. As shown in Table 5, our method exhibits a moderate inference time and relatively high memory usage. Specifically, TTST and DATSR require longer inference times due to the computational overhead of top-k attention mechanisms. EDiffSR and DiffBIR suffer from significantly reduced speed due to the iterative sampling nature of diffusion models. Regarding memory consumption, both our method and DiffBIR are built upon the pretrained Stable Diffusion model, resulting in a large number of parameters and increased memory usage. However, compared to the ControlNet-style architecture of DiffBIR , our adapter-based design is more lightweight, leading to lower memory consumption. The long inference time and massive memory consumption of SR methods based on naive Stable Diffusion models are common limitations. In the future, one-step distillation acceleration and model light-weighting techniques will be among our key directions for optimization.
@@ 6 Conclusion
In this paper, we proposed CRefDiff, a novel controllable reference-based diffusion framework for remote sensing image RefSR in real-world scenarios. To address the large resolution gap and complex land-cover variations commonly found in corss-sensor satellite imagery, CRefDiff integrates powerful generative priors from pretrained Stable Diffusion models with a dual-branch fusion mechanism that captures both local textures and global contextual information from the reference image. Specifically, a Change-Aware Attention block is designed to selectively inject fine-grained local details, while a Semantic Token Aggregation module distills global features from DiNOv2 embeddings to provide contextual guidance. Moreover, we introduce two optional inference strategies—Better Start for faster denoising and reference strength control for adaptive reference information modulation—making the framework both efficient and interactive.

### v2

@@ 5.6.5 Model Efficiency
To compare the runtime efficiency of different methods, we evaluate the inference time and GPU memory consumption for a $\times480$ image on an NVIDIA GeForce RTX 3090 GPU. As shown in Table 7, our method exhibits a moderate inference time and relatively high memory usage. Specifically, TTST and DATSR require longer inference times due to the computational overhead of top-k attention mechanisms. EDiffSR and DiffBIR suffer from significantly reduced speed due to the iterative sampling nature of diffusion models. Regarding memory consumption, both our method and DiffBIR are built upon the pretrained Stable Diffusion model, resulting in a large number of parameters and increased memory usage. However, compared to the ControlNet-style architecture of DiffBIR , our adapter-based design is more lightweight, leading to lower memory consumption. The long inference time and massive memory consumption of SR methods based on naive Stable Diffusion models are common limitations. In the future, one-step distillation acceleration and model light-weighting techniques will be among our key directions for optimization.
@@ 6 Conclusion
In this paper, we cast remote sensing RefSR as a multi-sensor data fusion task and propose CRefDiff, a novel controllable reference-guided diffusion framework designed for real-world remote sensing image RefSR. To address the large resolution gap and complex land-cover variations commonly found in corss-sensor satellite imagery, CRefDiff integrates powerful generative priors from a pretrained generative model with a dual-branch fusion mechanism that captures both local textures and global contextual information from the reference image. Specifically, a Change-Aware Attention block is designed to selectively inject fine-grained local details, while a Semantic Token Aggregation module distills global features from DiNOv2 embeddings to provide contextual guidance. Moreover, we introduce two optional inference strategies—Better Start for faster denoising and reference strength control for adaptive reference information modulation—making the framework both efficient and interactive.